# “It Felt Like I Was Left in the Dark”: Exploring Information Needs and Design Opportunities for Family Caregivers of Older Adult Patients in Critical Care Settings

### v2

@@ 1. Introduction
Building on these findings, we propose a set of design guidelines and implement an AI-based prototype system, CareMap, to support caregivers’ access to and understanding of ICU patient information. As illustrated in Figure 1, the system processes complex clinical information from the ICU and leverages three LLM-driven modules—summarization, extraction, and conversation—to generate caregiver-facing outputs, including daily summaries, treatment plans, and context-aware responses. These outputs are delivered through two core features: (1) a visual timeline of key medical events that streamlines access to systematic status updates, and (2) an LLM-based chatbot that enables personalized information inquiries through natural language conversation. We then conducted usability evaluations with 10 family caregivers to assess the prototype. Qualitative analysis revealed the potential of LLM-based systems to reduce caregivers’ cognitive burdens, improve their access to complex clinical information, and enhance engagement with care teams.
@@ 3.2. Interview Procedure
After each interview, the audio was transcribed. Two researchers then independently coded the transcripts using an open coding approach (Corbin and Strauss, 2014). They then discussed and reconciled their initial coding schemas, iteratively refining them for clarity and consistency. The finalized codebook (Appendix  2) was applied across all transcripts by both researchers, with any discrepancies resolved through team discussions until a consensus was reached.
@@ 5. CareMap: an AI-Based Prototype for Supporting Caregivers of Older Adult ICU Patients
Building on the design space in Section 4.3, we developed an AI-based prototype, CareMap, to help caregivers access and understand clinical information through two interconnected components: a visual timeline of patients’ medical events and an LLM-based chatbot to facilitate personalized information inquiries through conversations.
Figure 3. CareMap Prototype Designs: (A) A visual timeline that maps a patient’s medical events, featuring AI-generated summaries of lab test results, key treatment plans, and daily goals to provide a clear overview of the patient’s status. (B) An example of a context-aware LLM-based chatbot that provides insights into medical reports, helping users interpret test results and identify trends.
5.1. Visual Timeline of Patients’ Medical Events
@@
5.2. Context-Aware LLM-Based Chatbot
@@
Figure 4. Context-Aware LLM-Based Chatbot Prototype Design: (A) Fact Finder – Responds to user queries about ICU treatments and suggests relevant follow-up questions. (B) Data Interpreter – Explains lab test results and clinical data in an accessible manner. (C) Intention Prompter – Assists users in formulating meaningful questions for healthcare providers.
@@ 6.1. Participants and Recruitment
For the second stage of our study, we recruited the same ten caregivers (P6-P15) from Section 3 to conduct a usability evaluation of our prototype. Each prototype evaluation session was conducted remotely via Zoom and lasted approximately 30–45 minutes. To comply with usage regulations requiring approval for accessing MIMIC-III (Johnson et al., 2016) data, we integrated de-identified, synthesized patient data into our prototype. In this paper, we present only mock data to illustrate the user interface.
All sessions were recorded and transcribed with participant consent. For each design prototype, we allocated approximately 10–15 minutes for participants to explore its features. After a brief introduction to the interface, participants were encouraged to interact with the system, share their impressions, and evaluate its perceived usefulness in accessing and understanding a patient’s status. They also reflected on how they might have used the tool to obtain relevant medical information, identified the most helpful aspects, and suggested potential improvements to enhance overall usability. The detailed semi-structured interview protocol is provided in Appendix B. Following the interviews, two researchers independently analyzed the transcripts using an inductive approach (Thomas, 2006), identifying key insights related to usability and potential design improvements for each prototype, the final codebook is provided in Appendix 3.
@@ 6.2. Result
In the following section, we present the results of the usability evaluation, which captures valuable feedback from family caregivers on our prototype. Overall, participants greatly appreciated the prototype’s ability to enhance their access to clinical information by providing consistent, timely, and easy-to-understand updates. Meanwhile, they found the prototype’s capability to integrate with personalized patient EHRs particularly helpful for them to understand clinical information, as it could effectively respond to their specific inquiries, uncover additional relevant details, and offer deeper insights into the patient’s condition.
@@ 6.2.1. Visual Timeline of ICU Stay: Supporting Caregivers’ Access to Consistent Information
Overall, our participants found the visual timeline view to be a valuable tool for better supporting their access to dynamic patient status changes during ICU stays. Building on findings from the formative study (See Section 4.1.1), which highlighted caregivers’ difficulties with inconsistent information structuring and accessibility in ICU workflows, the timeline interface is designed to address this challenge. Caregivers often rely on fragmented communication with healthcare providers, making it difficult to piece together a comprehensive understanding of the patient’s condition and treatment process. A unified interface that uses a timeline to visualize patient progress in the ICU over time could address this gap by enabling them to track trends and identify critical events of the patient’s ICU stay. The timeline was useful not only for individual caregivers but also as a shared communication tool for the patient’s family. Participants noted that it helped bridge information gaps between family caregivers and other family members, making it easier to share updates on the patient’s condition.
@@
Additionally, the timeline helped reduce information overload — a common challenge in ICU environments — by organizing complex clinical data into a clear, time-based summary. This design minimized caregivers’ cognitive effort, allowing them to focus on understanding the patient’s overall status rather than manually piecing together fragmented updates. The interface supported on-demand access to critical information, enabling caregivers to check in as needed rather than constantly monitoring data:
@@
AI-generated summaries of lab test results effectively provided family caregivers with a quick overview of the patient’s status. However, participants emphasized the need to balance simplicity with sufficient medical context. While they found these summaries valuable for gaining an initial understanding—especially in high-stress ICU environments where cognitive load is a concern—they also expressed concerns about oversimplification. Reducing complex medical information to brief summaries risks omitting critical details, making it harder for caregivers to fully grasp the care process and provide effective support.
Since older adult patients often have multiple chronic conditions before being admitted to the ICU, caregivers prefer more than just AI-generated summaries of daily lab test results. Incorporating a patient’s prior medical history into these summaries is crucial, as it would help caregivers better understand acute conditions and how they relate to the patient’s overall health.
@@
For caregivers who cannot always be physically present in the ICU, access to detailed information about their loved one’s daily treatment plans and goals is crucial for maintaining a sense of connection and involvement in care. While our prototype outlines treatment plans—including key milestones, medications, dosages, and their purposes—to help caregivers stay engaged, many participants emphasized the need for even greater transparency to further enhance their involvement in the care process.
@@ 6.2.2. Context-Aware LLM-Based Chatbot: Enhancing Caregivers’ Understanding of ICU Information
While participants generally appreciated the chatbot, some raised concerns about verifying its responses and ensuring accuracy. For instance, P15 highlighted the importance of cross-checking AI-generated information with medical professionals, stating: “I wouldn’t like one hundred percent trust the chatbot that answers my question. But I think with time, I might verify [the] answers [by] speaking to a doctor.”
@@
Medical data in professional reports, often dense with jargon and numerical thresholds, can be overwhelming for caregivers unfamiliar with clinical term. Caregivers in our study expressed a need for clearer explanations of medical references to better understand their significance. The Data Interpreter helps address this challenge by providing definitions and contextual insights. For example, when faced with an elevated MCHC level, the chatbot can explain what the acronym stands for, describe its role in assessing blood health, and clarify what an abnormal result might indicate for the patient.
@@
Beyond providing reactive support, the Intention Prompter helps caregivers anticipate and navigate conversations with the clinical team more effectively. By suggesting questions based on the patient’s status, the system helps move conversations beyond repetitive exchanges, making them more meaningful and informative.
@@ 7.1. Beyond Ancillary Roles: Empowering Caregivers as Informal Healthcare Partners
In parallel, research in Computer-Supported Cooperative Work (CSCW) and Human-Computer Interaction (HCI) has adopted an ecological perspective to investigate the socio-technical ecosystems surrounding family caregivers in clinical settings. Empirical studies have systematically mapped caregivers’ multifaceted roles in clinical settings - from providing emotional support to coordinating appointments, transportation, and daily care during hospital stays (Miller et al., 2016; Pina et al., 2017; Barbarin et al., 2015). These findings collectively emphasize the importance of designing systems and practices that can better support caregivers’ evolving roles within complex clinical environments. Our research seeks to complement this body of work by further unpacking the complexities of caregivers’ informational needs as primary stakeholders, particularly within the high-stakes environment of the Intensive Care Unit (ICU). Specifically, we highlight the importance of two-way information flow and adaptive support systems that accommodate caregivers’ varying levels of health literacy. Our findings show that technology can empower caregivers to take on a more active role, not just as recipients of information, but as engaged participants in care coordination.
@@ 7.2.1. From the perspective of Human-AI Collaboration
Accessing and understanding clinical information is one of the most cognitively demanding tasks for family caregivers, especially in the high-pressure, fast-paced environment of the ICU. Based on findings from our two-stage study, we explored with participants how AI could help streamline their processes of accessing and making sense of information. A key direction for future design is to go beyond basic summarization or extraction. Instead, AI systems should engage in adaptive, ongoing interactions that learn from caregivers’ responses and progressively personalize information delivery.
In addition, AI systems deployed in clinical settings should be designed to foster appropriate user trust and expectations. Given the complexity of these systems and the high-stakes nature of ICU environments, it is critical to clearly communicate the system’s capabilities and limitations to caregivers. Caregivers in our study expressed concerns about the ambiguity of AI-generated clinical summaries, particularly when numerical data were presented without adequate explanation. We suggest that designers offer explanations tailored to the underlying technologies used in the system, such as clarifying how outputs were generated and presenting insights in concise, easy-to-understand formats.
@@ 7.2.2. From the perspective of CSCW
ICU caregiving is both collaborative and isolating—while families play a crucial role, they often feel excluded from clinical workflows structured around the ICU’s fast-paced environment. CSCW research explores ways to improve teamwork in healthcare by developing systems that help caregivers and healthcare providers share and interpret important information across different roles and data sources (Yang et al., 2024; Hao et al., 2024; Bowers et al., 2024). In our study, caregivers expressed a desire to supplement clinical data with their firsthand insights and their need for greater transparency and inclusion in coordinated care. To address this, future systems could allow caregivers to contribute their observations (e.g., concerns about specific treatments), with both clinicians and AI validating the information. This approach would facilitate structured information exchange and strengthen collaboration with healthcare providers. Additionally, if caregivers have the opportunity to provide input to the system and receive feedback from AI or healthcare providers, they can develop a better understanding of medical information as well as improve their health literacy.
7.3. Risks and Ethical Concerns of AI-based Clinical Information Technologies

### v3

@@ 1. Introduction
Building on these findings, we propose a set of design guidelines and implement an AI-based prototype system, CareMap, to support caregivers’ access to and understanding of ICU patient information. As illustrated in Figure 1, the system processes complex clinical information from the ICU and leverages three LLM-driven modules—summarization, extraction, and conversation—to generate caregiver-facing outputs, including daily summaries, treatment plans, and context-aware responses. These outputs are delivered through two core features: (1) a visual timeline of key medical events that streamlines access to systematic status updates, and (2) an LLM-based chatbot that enables personalized information inquiries through natural language conversation. We then conducted usability evaluations with 10 family caregivers to assess the prototype’s usability, and with 2 clinical experts to evaluate the quality of the system-generated outputs as well as provide usability feedback from a clinical perspective. Qualitative analysis revealed the potential of LLM-based systems to reduce caregivers’ cognitive burdens, improve their access to complex clinical information, and enhance engagement with care teams.
@@ 3.2. Interview Procedure
After each interview, the audio was transcribed. Two researchers then independently coded the transcripts using an open coding approach (Corbin and Strauss, 2014). They then discussed and reconciled their initial coding schemas, iteratively refining them for clarity and consistency. The finalized codebook (Appendix  3) was applied across all transcripts by both researchers, with any discrepancies resolved through team discussions until a consensus was reached.
@@ 5. CareMap: an AI-Based Prototype for Supporting Caregivers of Older Adult ICU Patients
Building on the design space outlined in Section 4.3, we developed an AI-based prototype system, CareMap, to support caregivers in accessing and understanding clinical information. An overview of the system architecture is shown in Figure 1. The protoype system is designed to seamlessly collect patient data from the EHR system, process it through the back-end structure (Section 5.2), and deliver the resulting information through the front-end interfaces to family caregivers (Section 5.1).
5.1. Front-end Interface Design
The front-end interface includes two interconnected components: a visual timeline of patients’ medical events and an LLM-based chatbot to facilitate personalized information inquiries through conversations.
Figure 3. CareMap Design: (A) A visual timeline that maps a patient’s medical events, featuring AI-generated summaries of lab test results, key treatment plans, and daily goals to provide a clear overview of the patient’s status. (B) An example of a context-aware LLM-based chatbot that provides insights into medical reports, helping users interpret test results and identify trends.
5.1.1. Visual Timeline of Patients’ Medical Events
@@
5.1.2. Context-Aware LLM-Based Chatbot
@@
Figure 4. Context-Aware LLM-Based Chatbot Design: (A) Fact Finder – Responds to user queries about ICU treatments and suggests relevant follow-up questions. (B) Data Interpreter – Explains lab test results and clinical data in an accessible manner. (C) Intention Prompter – Assists users in formulating meaningful questions for healthcare providers.
5.2. Backend Structure Design
The back-end architecture of CareMap stores, processes, and analyzes clinical data collected from Electronic Health Records (EHRs) to support the two front-end components described in Section 5.1. It consists of three core modules: the Data Extraction Module, the Summarization Module, and the Conversational Module.
Our back-end algorithms are inspired by prior work on patient-centered EHR interaction (Zhang et al., 2023) and longitudinal clinical reasoning using LLMs (Kruse et al., 2025). For example, Kruse et al. (2025) systematically evaluated state-of-the-art LLMs on multi-day Assessment & Plan generation and discharge summarization using MIMIC-III data, benchmarking three prompting strategies—direct generation, Retrieval-Augmented Generation (RAG) (Gao et al., 2023; Lewis et al., 2020), and Chain-of-Thought (CoT) (Wei et al., 2022). Inspired by this approach, our Data Extraction Module leverages LLM-based processing to standardize and structure EHR data (e.g., lab tests, medications, vital signs) and temporally order clinical events to support retrieval and generation, while our Summarization Module uses LLMs to retrieve multi-day evidence and perform temporal reasoning across structured and unstructured data to produce daily summaries.
Complementary to this, Zhang et al. (2023) developed the NoteAid EHR Interaction Pipeline, which leverages LLMs — specifically GPT-3.5-Turbo and GPT-4 — to construct a synthetic dialogue dataset for patient education on EHR notes. Their system pairs a Mock-Patient LLM agent and an Assistant LLM agent to simulate multi-round conversations (three rounds each) that perform Q&A and explanation tasks on unstructured discharge notes from MIMIC-III (Johnson et al., 2016) and MADE (Jagannatha et al., 2019). This approach yielded 43,504 synthetic interaction instances, demonstrating the potential ofLLMs to enhance patients’ understanding of their medical records . Building on this paradigm, our Conversational Module similarly grounds its responses in patient charts and produces plain-language explanations tailored for caregivers.
To further assess the clinical validity of the AI-generated outputs, we consulted two experienced board-certified clinicians to conduct an in-depth evaluation. We selected four pairs of AI-generated summaries and their corresponding EHR inputs from Zhang et al. (2023)’s example. The clinicians rated each summary on five criteria—Overall Accuracy, Hallucination, Readability, Specificity, and Overall Quality. The detailed questionnaire used for this evaluation is provided in Appendix B.2.
The results of this evaluation are shown in Figure 5. As illustrated, the clinicians gave consistently high ratings for the overall quality of the AI-generated outputs.(M = 4), including Overall Accuracy (M = 4), Hallucination (M = 4.5), Readability (M = 4.5), and Specificity (M = 5).
Figure 5. Evaluation scores given by two clinician experts (C1 and C2) across five criteria—Overall Accuracy, Hallucination, Readability, Specificity, and Overall Quality—for the AI-generated outputs.
@@ 6.1. Participants and Recruitment
For the second stage of our study, we recruited the same ten caregivers (P6-P15) from Section 3 to conduct a usability evaluation of our prototype. We also recruited 2 clinicians with significant experience in this area by convenience sampling (Etikan et al., 2016), as detailed in table 2. Recruitment involved leveraging professional networks and referrals from colleagues. Each prototype evaluation session was conducted remotely via Zoom and lasted approximately 30–45 minutes. To comply with usage regulations requiring approval for accessing MIMIC-III (Johnson et al., 2016) data, we integrated de-identified, synthesized patient data into our prototype. In this paper, we present only mock data to illustrate the user interface.
C# Gender Department Job Title Year of Practice
C1 Female Critical Care Physician 13 Years
C2 Female Emergency Medicine Physician 10 Years
Table 2. Clinician experts’ background summary.
All sessions were recorded and transcribed with participant consent. For each design prototype, we allocated approximately 10–15 minutes for family caregivers to explore its features. After a brief introduction to the interface, caregivers were encouraged to interact with the system, share their impressions, and evaluate its perceived usefulness in accessing and understanding a patient’s status. They also reflected on how they might have used the tool to obtain relevant medical information, identified the most helpful aspects, and suggested potential improvements to enhance overall usability. In parallel, two clinical experts reviewed the same prototypes. They were asked to examine the AI-generated outputs (daily summaries, treatment plans, and chatbot responses) for Overall Accuracy, Hallucination, Readability, Specificity, and Overall Quality (as detailed in Section 5.2), and to provide usability feedback on the interface from a clinical perspective.
The detailed semi-structured interview protocols are provided in Appendix B. Following the interviews, two researchers independently analyzed the transcripts using an inductive approach (Thomas, 2006), identifying key insights related to usability and potential design improvements for each prototype; the final codebook is provided in Appendix 4.
@@ 6.2. Result
In the following section, we present the results of the usability evaluation, which captures valuable feedback from family caregivers and clinician experts on our prototype. Overall, family caregivers greatly appreciated the prototype’s ability to enhance their access to clinical information by providing consistent, timely, and easy-to-understand updates. Meanwhile, they found the prototype’s capability to integrate with personalized patient EHRs particularly helpful for them to understand clinical information, as it could effectively respond to their specific inquiries, uncover additional relevant details, and offer deeper insights into the patient’s condition. In parallel, clinicians provided positive feedback on its potential to support accurate information delivery and to help caregivers interpret complex medical data in a patient-specific context.
@@ 6.2.1. Visual Timeline of ICU Stay: Supporting Caregivers’ Access to Consistent Information
Overall, caregivers found the visual timeline view to be a valuable tool for better supporting their access to dynamic patient status changes during ICU stays. Building on findings from the formative study (See Section 4.1.1), which highlighted caregivers’ difficulties with inconsistent information structuring and accessibility in ICU workflows, the timeline interface is designed to address this challenge. Caregivers often rely on fragmented communication with healthcare providers, making it difficult to piece together a comprehensive understanding of the patient’s condition and treatment process. A unified interface that uses a timeline to visualize patient progress in the ICU over time could address this gap by enabling them to track trends and identify critical events of the patient’s ICU stay. also provided comments about its potential to help families recognize the passage of time during the ICU stay.
“ (this) might provide some relevant time to the family,… because I think sometimes the families don’t realize how long it has been.”(C1)
The timeline was useful not only for individual caregivers but also as a shared communication tool for the patient’s family. Some caregivers noted that it helped bridge information gaps between family caregivers and other family members, making it easier to share updates on the patient’s condition.
@@
Additionally, the timeline helped reduce information overload — a common challenge in ICU environments — by organizing complex clinical data into a clear, time-based summary. As patients often undergo many procedures each day, as commented by clinician:
“My patients have many procedures, and it’s really nice to have a tab that says: on the first day after the operating room, they had a central line placed. (On) the second day, they had a tracheostomy. Those things are actually quite nice to [be] organized.”(C2)
This design minimized caregivers’ cognitive effort, allowing family caregivers to focus on understanding the patient’s overall status rather than manually piecing together fragmented updates. The interface supported on-demand access to critical information, enabling caregivers to check in as needed rather than constantly monitoring data:
@@
AI-generated summaries of lab test results effectively provided family caregivers with a quick overview of the patient’s status. However, both clinicians and caregivers emphasized the need to balance simplicity with sufficient medical context. While summaries were found valuable for gaining an initial understanding—especially in high-stress ICU environments where cognitive load is a concern—they also expressed concerns about oversimplification. Reducing complex medical information to brief summaries risks omitting critical details, making it harder for caregivers to fully grasp the care process and provide effective support. Since older adult patients often have multiple chronic conditions before being admitted to the ICU, caregivers prefer more than just AI-generated summaries of daily lab test results. Incorporating a patient’s prior medical history into these summaries is crucial, as it would help caregivers better understand acute conditions and how they relate to the patient’s overall health.
@@
For caregivers who cannot always be physically present in the ICU, access to detailed information about their loved one’s daily treatment plans and goals is crucial for maintaining a sense of connection and involvement in care. While our prototype outlines treatment plans—including key milestones, medications, dosages, and their purposes—to help caregivers stay engaged, many participants emphasized the need for even greater transparency to further enhance their involvement in the care process. Moreover, as mentioned in Section 5.2, clinicians expressed slight concerns about readability (M = 4.5). Specifically, they emphasized the need to clarify which treatment plans should be prioritized when presented to caregivers:
… [the treatment plan] gonna be really challenging to be put into a summary for family caregivers, most of these patients are on ten plus medications..(C2)
@@ 6.2.2. Context-Aware LLM-Based Chatbot: Enhancing Caregivers’ Understanding of ICU Information
While caregivers generally appreciated the chatbot, some raised concerns about verifying its responses and ensuring accuracy. For instance, P15 highlighted the importance of cross-checking AI-generated information with medical professionals, stating: “I wouldn’t like one hundred percent trust the chatbot that answers my question. But I think with time, I might verify [the] answers [by] speaking to a doctor.”
@@
Medical data in professional reports, often dense with jargon and numerical thresholds, can be overwhelming for caregivers unfamiliar with clinical term. Caregivers in our study expressed a need for clearer explanations of medical references to better understand their significance. The Data Interpreter helps address this challenge by providing definitions and contextual insights. Clinicians appreciated its ability to help caregivers address complex questions about lab test results, while expressing slight concerns about potential hallucinations in numerical interpretation (M = 4.5), as mentioned in Section 5.2:
“.. because we’ll be in the middle of rounds, and families will ask us very specific questions (like) I can see my family members’ results constantly, and they’re asking about every little number, including numbers I don’t even look at, (So I tell them) it doesn’t matter…”(C1)
@@
Beyond providing reactive support, the Intention Prompter helps caregivers anticipate and navigate conversations with the clinical team more effectively. Clinicians also emphasized that these questions should be tailored based on the patient’s specific condition: “…if they brought it to me with the context of their own ideas and interpretation of what’s been going on, it could be helpful…”(C2). By suggesting questions based on the patient’s status, the system helps move conversations beyond repetitive exchanges, making them more meaningful and informative.
@@ 7.1. Beyond Ancillary Roles: Empowering Caregivers as Informal Healthcare Partners
In parallel, research in Computer-Supported Cooperative Work (CSCW) and Human-Computer Interaction (HCI) has adopted an ecological perspective to investigate the socio-technical ecosystems surrounding family caregivers in clinical settings. Empirical studies have systematically mapped caregivers’ multifaceted roles in clinical settings - from providing emotional support to coordinating appointments, transportation, and daily care during hospital stays (Miller et al., 2016; Pina et al., 2017; Barbarin et al., 2015). These findings collectively emphasize the importance of designing systems and practices that can better support caregivers’ evolving roles within complex clinical environments. Our research seeks to complement this body of work by further unpacking the complexities of caregivers’ informational needs as primary stakeholders, particularly within the high-stakes environment of the Intensive Care Unit (ICU). Moreover, with the add-on feedback from clinicians, they recognized the challenges faced by caregivers and how our system can help bridge informational gaps and support their understanding of patients’ conditions. Specifically, we highlight the importance of two-way information flow and adaptive support systems that accommodate caregivers’ varying levels of health literacy. Our findings show that technology can empower caregivers to take on a more active role, not just as recipients of information, but as engaged participants in care coordination.
@@ 7.2.1. From the perspective of Human-AI Collaboration
Accessing and understanding clinical information is one of the most cognitively demanding tasks for family caregivers, especially in the high-pressure, fast-paced environment of the ICU. Based on findings from our two-stage study, we explored with caregivers how AI could help streamline their processes of accessing and making sense of information. A key direction for future design is to go beyond basic summarization or extraction. Instead, AI systems should engage in adaptive, ongoing interactions that learn from caregivers’ responses and progressively personalize information delivery.
In addition, AI systems deployed in clinical settings should be designed to foster appropriate user trust and expectations. Given the complexity of these systems and the high-stakes nature of ICU environments, it is critical to clearly communicate the system’s capabilities and limitations to caregivers. While clinicians generally gave high overall ratings for the AI-generated outputs, caregivers in our study expressed concerns about the ambiguity of these summaries, particularly when numerical data were presented without adequate explanation. We suggest that designers offer explanations tailored to the underlying technologies used in the system, such as clarifying how outputs were generated and presenting insights in concise, easy-to-understand formats.
@@ 7.2.2. From the perspective of CSCW
ICU caregiving is both collaborative and isolating—while families play a crucial role, they often feel excluded from clinical workflows structured around the ICU’s fast-paced environment. CSCW research explores ways to improve teamwork in healthcare by developing systems that help caregivers and healthcare providers share and interpret important information across different roles and data sources (Yang et al., 2024; Hao et al., 2024; Bowers et al., 2024). In our study, caregivers expressed a desire to supplement clinical data with their firsthand insights and their need for greater transparency and inclusion in coordinated care. To address this, future systems could allow caregivers to contribute their observations (e.g., concerns about specific treatments), with both clinicians and AI validating the information. This approach would facilitate structured information exchange and strengthen collaboration with healthcare providers. If caregivers have the opportunity to provide input to the system and receive feedback from AI or healthcare providers, they can develop a better understanding of medical information as well as improve their health literacy. Furthermore, as clinicians intend to use this interface to update patient information, it might bridge the gap between clinical decision-making and caregivers’ understanding, fostering more aligned expectations and collaborative care.
7.3. Risks, Technical and Ethical Concerns of AI-based Clinical Information Technologies